OpenStudy (anonymous):

PLZZZ!! HELP WITH MATH!!! I'LL GIVE MEDALS, FANS AND TESTIMONIES :) A cone has a volume of 5 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of? 15 in3 20 in3 25 in3 30 in3

OpenStudy (anonymous):

@linaboo123

OpenStudy (anonymous):

Volume of Cone: 1/3 Bh where B is the area of the base 1/3 Bh = 5 Bh = 15 B = 15/h The cylinder has the same base area as the cone. Volume of Cylinder: V = B*h B = V/h and B = 15/h The height of the cone and the cylinder are the same. V/h = 15/h Volume of Cylinder = 15.

OpenStudy (anonymous):

Its b

OpenStudy (anonymous):

Its A.

OpenStudy (anonymous):

@Dredge u know its b

OpenStudy (anonymous):

lol @linaboo123 hows it b tho??

OpenStudy (anonymous):

i have a smart math teacher

OpenStudy (anonymous):

It's A. -.-

OpenStudy (anonymous):

@SmokeyRiver that v=1/3bh is the formula for the volume of a pyramid, why do we use it here?

OpenStudy (anonymous):

@linaboo123 might be right

OpenStudy (anonymous):

i am